OpenStudy (anonymous):

Write the quadratic function in the form. f(x)=-2x^2-8x-12

OpenStudy (anonymous):

@liz_mustang @Librarian @triciaal @e.mccormick @Compassionate @milkacha @fateal

OpenStudy (anonymous):

@ganeshie8

OpenStudy (fateal):

like this???

OpenStudy (anonymous):

Write the quadratic function in the form f(x)=a(x-h)^2+k

OpenStudy (anonymous):

use completing the square formula/steps

OpenStudy (anonymous):

i don't know how to do that

OpenStudy (anonymous):

well first of all try to factor out something that goes evenly in the 3 terms

OpenStudy (anonymous):

oh i actually do!!! Thats pre cal stuff, complete the square!

OpenStudy (anonymous):

give it a try before we give you answer, have you guys done any examples in class?

OpenStudy (anonymous):

@liz_mustang can you help me? @juniro no i in online classes

OpenStudy (anonymous):

ok first factor out -2 from the equation

OpenStudy (anonymous):

this is how product should look after factoring

OpenStudy (anonymous):

then you can complete square

OpenStudy (anonymous):

let me check if that correct

OpenStudy (anonymous):

thats not the answer yet, you need to complete the square after you factor out -2

OpenStudy (anonymous):

this is example of how to complete the square

OpenStudy (triciaal):

the posted question is not clear.

OpenStudy (anonymous):

you would use (b/2)^2 b=the number next to the x variable

OpenStudy (anonymous):

Write the quadratic function in the form. <<<f(x)=a(x-h)^2+k f(x)=-2x^2-8x-12

OpenStudy (anonymous):

does the eponent 2 are added to -8x

OpenStudy (triciaal):

no

OpenStudy (anonymous):

use the distribution properties

OpenStudy (anonymous):

so where does the exponent goes ?

OpenStudy (triciaal):

standard a = -2 b = -8 and c = -12 h = -b/2a and k = c + h^2

OpenStudy (anonymous):

what did you get when you factor out -2? and after use (b/2)^2?

OpenStudy (anonymous):

what x? @triciaal

OpenStudy (triciaal):

correct

OpenStudy (triciaal):

@sunnycali what do you mean "what x ? "